OpenStudy (anonymous):

solve: 2x+5/2x+3=x+3/x-1

OpenStudy (anonymous):

(2x+5)*(x-1) = (x+3)(2x+3) 2x^2 + 3x -5 = 2x^2 +9x+9 -6x = 14 x = -14/6

OpenStudy (anonymous):

there r only 3 anwers -2 1/3, 1/3, () please do not give up

OpenStudy (anonymous):

No i am absolutely sure the answer i gave is right. Or the question is presented in a wrong way.

OpenStudy (anonymous):

i know this online school is dumb the problem states: solve" 2x+5/2x+3 = x+3/x-1

OpenStudy (anonymous):

The the answer is -7/3 for sure :) There is no way you get -2 or 1/3 from that.

OpenStudy (ybarrap):

Note that -2 1/3 = -7/3.